# Wave Attenuation in Drifting Sea Ice: A Mechanistic Model for Observed Decay Profiles

### v1

@@ Abstract
Wave-sea ice interactions shape the transition zone between open ocean and pack ice in the polar regions. Most theoretical paradigms, implemented in coupled wave-sea ice models, predict exponential decay of the wave energy but recent observations deviate from this behaviour. Expanding on a framework based on wave energy dissipation due to ice-water drag, we account for drifting sea ice to derive an improved model for wave energy attenuation. Analytical solutions replicate the observed non-exponential wave energy decay and the spatial evolution of the effective attenuation rate in Antarctic sea ice.
@@ 1 Introduction
The interaction of Southern Ocean waves propagating into the sea ice for hundreds of kilometres is a defining feature of the Antarctic MIZ (Kohout et al., 2014; Stopa et al., 2018; Brouwer et al., 2022; Day et al., 2024). Under persistent waves (Derkani et al., 2020), the MIZ comprises, for the most part, unconsolidated floes (<10<10 m in diameter) in high concentration ($80\%$; Alberello et al., 2019; Day et al., 2024). In these conditions, with little internal stresses (Alberello et al., 2020; Womack et al., 2022), wind is the main driver of sea ice drift (Holland & Kwok, 2012; Kwok et al., 2017). Through air-ice drag coefficients of order $10^{-3}$ (Wamser & Martinson, 1993; Elvidge et al., 2016), sea ice drifts at 0.1–0.4 m/s with low variability over large scales (10–100 km; Holland & Kwok, 2012).
@@
Wave forcing decays with distance into the MIZ, with the attenuation rate assumed to be proportional to the wave frequency (e.g. Meylan et al., 2018). The exact frequency dependency varies according to the dominant attenuation mechanism, and various options are implemented in operational third generation spectral wave models (e.g. WaveWatchIII; Montiel et al., 2025) that are being coupled into global- and hemispheric-scale sea ice models. Most models predict a constant attenuation rate in homogeneous sea ice conditions and, consequently, exponential wave energy decay.
@@
Overall, measurements conform reasonably well to the exponential attenuation paradigm and agree on the magnitude of the attenuation rate $\alpha$, within the range ($10^{-4}$ $^{-1}<\alpha<10^{-5} m-$1; Kohout et al., 2020; Montiel et al., 2022). However, recent satellite measurements of wave height in the Antarctic MIZ across seasons revealed an almost linear increase of the average attenuation rate with distance from the sea ice edge (up to $\approx 10\times$; Voermans et al., 2025) with individual transects showing sharp peaks in attenuation rate. Possible causes of deviation from the exponential profile were attributed to the misalignment between the transect and wave direction, and inhomogeneous sea ice conditions (sea ice concentration and thickness). Lacking a firm theoretical foundation, Voermans et al. (2025) propose an empirical wave attenuation formulation as an interim solution for implementation in wave models.
@@
Amongst the various theoretical wave attenuation models, the ones that attribute wave energy dissipation to quadratic ice-water drag deviate from exponential decay in form. Ice-water drag models capture the bulk of the energy loss when sea ice comprises floes with small diameter, relative to wavelength, in high concentration (Shen & Squire, 1998; Kohout et al., 2011; Herman et al., 2019), this makes them well suited to the Antarctic MIZ. The attenuation rate due to drag is proportional to the square of the relative velocity between fluid orbital motion and sea ice (Herman et al., 2019). Assuming a stationary a semi-infinite sea ice domain, analytical, non-exponential, solutions for the amplitude profile have been obtained under slightly different approaches by Kohout et al. (2011), for the amplitude envelope, and Herman et al. (2019), for the phase-averaged amplitude. The former is a lower bound for the amplitude, i.e. higher bound for the attenuation rate.
Here, we develop an analytical framework for the attenuation of monochromatic waves due to drag, explicitly accounting for constant sea ice drift. We extend the previous works of Kohout et al. (2011) and Herman et al. (2019) where drift was neglected on the grounds that, in typical oceanic conditions, drift velocities are two orders of magnitude smaller than the wave group velocity. Despite its apparently negligible effect, we reveal distinctive consequences of including such drift, most notably the emergence of a location into the ice field where the wave amplitude vanishes and the attenuation rate becomes unbounded. We further demonstrate the capability of this enhanced model to reproduce recent observations of wave attenuation in the Antarctic MIZ.
@@ 2 Mathematical Formulation and Analytical Solutions
where $c_{g}$ denotes the group velocity and, similarly to Kohout et al. (2011), the effect of the ice $(S_{\text{ice}}$) is explicitly expressed as a linear superposition of all the mechanisms with constant decay, that lead to exponential attenuation ($S_{\text{exp}}$), and the ones due to ocean-ice skin drag drag $(S_{\text{sd}}$). Here, in agreement with most wave-in-ice parametrisations in WaveWatchIII (Montiel et al., 2025), we assume the open water dispersion relation between angular frequency $\omega$ and wave-number $k$ to hold in sea ice.
The transport equation in terms of wave amplitude $a(x,t)$, using $E\propto\rho ga^{2}$ where $\rho$ is water density and $g$ acceleration due to gravity, becomes:
@@
where $\alpha_{\textrm{exp}}$ is the constant attenuation rate for all other contributing non skin-drag mechanisms. Its inclusion provides flexibility for future operational model implementations.
In ice covered regions, for a monochromatic wave, the rate of change of energy per unit area due to skin drag is proportional to $|u_{\text{orb}}-v|^{3}$ (Shen & Squire, 1998), where $u_{\text{orb}}=U\sin{(\phi)}$ is the wave orbital velocity at the ice-water interface, here evaluated at the free surface under the hypothesis of thin sea ice, and $v$ is the ice drift velocity. The magnitude of the orbital velocity is $U=a\Omega$, where $\Omega=gk/\omega$, and $\phi$ is time-dependent phase. Phase averaging, we obtain:
$s_{\text{sd}}=-\frac{C_{d}}{g}\frac{1}{2\pi}\int_{0}^{2\pi}\left|a\Omega\sin(\phi)-v\right|^{3}d\phi,$ (4)
with $C_{d}$ the drag coefficient, assumed constant in homogeneous sea ice conditions.
@@ 2.1 Solving the Amplitude Profile
Equation  5 represents a challenging nonlinear ordinary differential equation. Also note that a polynomial solution in $a$ for the integral of the skin drag (Eq. 4) exists only for $|v|\geq a\Omega$. In effort to enable closed form analytical solutions for Eq. 5, we consider an asymptotic expression for the skin drag in the regime $|v|\ll a\Omega$, then extended into the region $|v|<a\Omega$, therefore obtaining:
@@
Note that, in oceanic conditions, $c_{g}$ is two orders of magnitude lager than the wind and current induced drift, i.e. $\mathcal{O}(10)\,\mathrm{m/s}$ versus $\mathcal{O}(0.1)\,\mathrm{m/s}$, therefore $|v|\ll c_{g}$ implying that, as the results of Eq. 7 are strictly positive, the amplitude $a$ monotonically decreases in Eq. 5. For $a_{0}\Omega>|v|$, we define $\hat{x}^{*}$ to be the location into the ice domain such that $a(\hat{x}^{*})\Omega=|v|$. Therefore, the condition $|v|\lessgtr a\Omega$ corresponds to $\hat{x}\lessgtr\hat{x}^{*}$ in the physical domain. In contrast, for $a_{0}\Omega\leq|v|$, one obtains $\hat{x}^{*}=0$.
@@
where the piecewise solution components result in:
@@
which controls the roots of the quadratic polynomial in $a$ affecting the solution in the $\hat{x}<\hat{x}^{*}$ region. Note that the physical interpretation of $\delta$ will be discussed in §2.2.
The solution has a few interesting features. It is nonlinear, i.e. drag effects and other mechanisms cannot be added as linear superposition. The solutions for $\hat{x}<\hat{x}^{*}$ in Eq. 9 differs in form but possess similar qualitative behaviour, where we note that $\hat{x}^{*}$ depends on $\delta$. For a wave with initial orbital velocity magnitude $U_{0}<|v|$ we obtain a single branched solution, i.e. $a_{B}$.
@@
such that $a_{B}(\hat{x}=\hat{x}_{\text{end}})=0$. The extinction location is unique to dissipation due to drift induced skin drag and defines the extent of the wave affected region. It moves closer to the sea ice edge for increasing $C_{D}$ coefficient and $\alpha$ (see solution in red, for $v=0.05$ m/s and $\alpha=7\times 10^{-6}$ m/s $\delta=3.4$], vs in blue, for $v=0.05$ m/s and $\alpha=0$ $\delta=\infty$], in Fig. 1a), while its dependence on drift is not trivial. Such extinction locations are not present in the solutions of Kohout et al. (2011) and Herman et al. (2019), shown dashed green and orange respectively, for which $\alpha=0$ and $v=0$.
@@
Figure 1: Sample amplitude (a) and corresponding attenuation (b) profiles according to our model and formulations by Kohout et al. (2011) and Herman et al. (2019). Dashed lines denote $x^{*}$. Amplitude is normalised by the initial amplitude and distance by $x_{\text{end}}$ for $\alpha=0$. Attenuation rate is normalised by $\alpha$.
Fig. 1a showcases the effect of drift. The attenuation is stronger when drift is present, see solution in red vs in purple ($v=0$ m/s and $\alpha=7\times 10^{-6}$). The increased attenuation due to other mechanisms also manifests when drift is present, see solutions in blue vs in red.
@@ 2.2 Predicting the Attenuation Rate
To allow better comparison with results as they appear in the literature, we compute the effective attenuation rate in the moving frame in the ice covered region defined as
@@
with the global minimum located in the region $\hat{x}<\hat{x}^{*}$ if $a_{0}\Omega\geq 3\sqrt{2}|v|/2$, whereas for lower orbital velocity at the sea ice edge the attenuation monotonically increases. This result allows for a physical interpretation of the dimensionless parameter $\delta$: for $\delta>1$ wave attenuation is dominated by drift induced skin drag and for $\delta<1$ all other mechanisms dominate.
For drifting sea ice, the attenuation rate grows unbounded in the neighbourhood of the extinction location (see solutions in red and blue in Fig. 1b), i.e.  $\alpha_{\text{eff}}(\hat{x})\sim[2(x_{\text{end}}-\hat{x})]^{-1}$, therefore the extinction location affects the magnitude of the attenuation rate in this region. When drift is absent, $\hat{x}^{*}\rightarrow\infty$ and the solution for $\hat{x}<\hat{x}^{*}$ applies to the entire domain. Skin drag dominates attenuation close to the sea ice edge whereas other processes dominate deeper into the sea ice, where we retrieve the classical exponential attenuation, i.e. $\alpha_{\text{eff}}\rightarrow\alpha$ for $\hat{x}\rightarrow\infty$ (see solution in purple in Fig. 1b; the grey dashed line denotes $\alpha_{\text{eff}}=\alpha$).

### v2

@@ Abstract
Wave-sea ice interactions shape the transition zone between open ocean and pack ice in the polar regions. Most theoretical paradigms, implemented in coupled wave-sea ice models, predict exponential decay of the wave energy but some recent observations deviate from this behaviour. Expanding on a framework based on wave energy dissipation due to ice-water drag, we account for drifting sea ice to derive an improved model for wave energy attenuation. Analytical solutions replicate the observed non-exponential wave energy decay and the spatial evolution of the effective attenuation rate in Antarctic sea ice.
@@ 1 Introduction
The interaction of ocean waves propagating into the sea ice is a defining feature of the Arctic and Antarctic MIZ (Kohout et al., 2014; Stopa et al., 2018; Alberello et al., 2022; Brouwer et al., 2022; Day et al., 2024), and in the Southern Ocean this region extends for hundreds of kilometres. Wave energy decays with distance into the MIZ, with the attenuation rate $\alpha$, i.e. $E(x)\propto\exp(-\alpha x)$, assumed to be proportional to the wave frequency $f$ to some power ($\alpha\propto f^{p}$; e.g. Meylan et al., 2018). The exact frequency dependency varies according to the dominant attenuation mechanism, and various options are implemented in operational third generation spectral wave models (e.g. WaveWatchIII; Montiel et al., 2025) that are being coupled into global- and hemispheric-scale ocean-sea ice models. Most models predict a constant attenuation rate in homogeneous sea ice conditions and, consequently, exponential wave energy decay.
@@
Overall, measurements conform reasonably well to the exponential attenuation paradigm and agree on the magnitude of the attenuation rate $\alpha$, within the range ($10^{-4}$ $^{-1}<\alpha<10^{-5} m-$1; Kohout et al., 2020; Montiel et al., 2022). However, recent satellite measurements of wave height in the Antarctic MIZ across seasons revealed an almost linear increase of the average attenuation rate with distance from the sea ice edge (up to $\approx 10\times$; Voermans et al., 2025) with individual transects showing sharp peaks in attenuation rate. Possible causes of deviation from the exponential profile were attributed to the misalignment between the transect and wave direction, and inhomogeneous sea ice conditions (sea ice concentration and thickness). Lacking a firm theoretical foundation, Voermans et al. (2025) propose an empirical wave attenuation formulation as an interim solution for implementation in wave models, i.e. $\alpha=\beta x/x_{\text{MIZ}}-\gamma$ in which $\beta$ and $\gamma$ are empirical fitting parameters and $x_{\text{MIZ}}$ is the extent of the MIZ. The formulation leads to exponential energy decay at a rate $\beta x^{2}$.
@@
Amongst the various theoretical wave attenuation models, the ones that attribute wave energy dissipation to quadratic ice-water drag deviate from exponential decay in form. Ice-water drag models capture the bulk of the energy loss when sea ice comprises floes with small diameter, relative to wavelength, in high concentration (Shen & Squire, 1998; Kohout et al., 2011; Herman et al., 2019) as further supported by experimental observations (Smith & Thomson, 2020). This makes these models well suited to the Antarctic MIZ. The attenuation rate due to drag is proportional to the square of the relative velocity between fluid orbital motion and sea ice (Herman et al., 2019). Assuming a stationary semi-infinite sea ice domain, analytical, non-exponential, solutions for the amplitude profile have been obtained under slightly different approaches by Kohout et al. (2011), for the amplitude envelope, and Herman et al. (2019), for the phase-averaged amplitude. The former is a lower bound for the amplitude, i.e. a higher bound for the attenuation rate.
The drift of sea ice is a feature of MIZ dynamics throughout the year (Kwok et al., 2017; Day et al., 2024). The Southern Ocean, under persistent waves (Derkani et al., 2020), is characterised by unconsolidated small floes ($\approx 10$ m in diameter) in high concentration ($80\%$; Alberello et al., 2019; Day et al., 2024) with little internal stresses (Alberello et al., 2020; Womack et al., 2022). In these conditions, wind is the main driver of sea ice drift (Holland & Kwok, 2012; Kwok et al., 2017). Through air-ice drag coefficients of order $10^{-3}$ (Wamser & Martinson, 1993; Elvidge et al., 2016), sea ice drifts at 0.1–0.4 m/s with low variability over large scales (10–100 km; Holland & Kwok, 2012). Nevertheless, the drift of the MIZ is particularly intense during extreme events when concurrent energetic ocean wave are able to penetrate hundreds of kilometres into the sea ice, as demonstrated by recent observations (Vichi et al., 2019; Alberello et al., 2020; Womack et al., 2022, 2024). These events are likely the main contributors to the overall MIZ dynamics. However, to the best of our knowledge, the interplay between drift and wave attenuation has not been explored before.
Several of the world’s leading meteorological agencies have moved to using coupled atmosphere–ocean–sea-ice forecasting systems in the last few years to improve forecasting of near-surface temperature, humidity, winds and low-level clouds, as well as sea-ice distribution in the polar regions. In this regard, prediction in the MIZ is proving a particularly difficult challenge (Day et al., 2022; Barrell et al., 2025). One aspect of focussed development is the coupling of physics-based ocean wave and sea-ice models, within a dynamical ocean model, so subject to drift (Kousal et al., 2026), making this study timely with the potential to inform this development.
Here, we develop a new analytical framework for the attenuation of monochromatic waves due to drag, explicitly accounting for constant sea ice drift.
We extend the previous works of Kohout et al. (2011) and Herman et al. (2019) where drift was neglected on the grounds that, in typical oceanic conditions, drift velocities are two orders of magnitude smaller than the wave group velocity.
Despite its apparently negligible effect, we reveal distinctive consequences of including such drift, most notably the emergence of a location into the ice field where the wave amplitude vanishes and the attenuation rate becomes unbounded.
We further demonstrate the capability of this enhanced model to reproduce recent observations of wave attenuation in the Antarctic MIZ.
@@ 2 Mathematical Formulation and Analytical Solutions
where $c_{g}$ denotes the group velocity
and, similarly to Kohout et al. (2011), the effect of the ice $(S_{\text{ice}}$) is explicitly expressed as a linear superposition of all the mechanisms with constant decay, that lead to exponential attenuation ($S_{\text{exp}}$), and the ones due to ocean-ice skin drag $(S_{\text{sd}}$). Note that the assumption of exponential decay for all the mechanisms other than drag will facilitate analytic handling. Here, in agreement with most wave-in-ice parametrisations in WaveWatchIII (Montiel et al., 2025), we assume the open water dispersion relation between angular frequency $\omega$ and wave-number $k$ to hold in sea ice.
The transport equation in terms of wave amplitude $a(x,t)$, using $E=\frac{1}{2}\rho ga^{2}$ where $\rho$ is water density and $g$ acceleration due to gravity, becomes:
@@
where $\alpha_{\textrm{exp}}$ is the constant amplitude attenuation rate for all other contributing non skin-drag mechanisms e.g. wave scattering (Wadhams, 1978), viscous dissipation (Weber, 1987; Keller, 1998) and visco-elasticity (Robinson & Palmer, 1990; Fox & Squire, 1994; Wang & Shen, 2010). Its inclusion provides flexibility for future operational model implementations.
In ice covered regions, for a monochromatic wave, the skin drag stress at the ice-ocean interface follows a standard quadratic law, $\tau=\rho C_{d}|u_{\text{orb}}-v|(u_{\text{orb}}-v)$ (Shen & Squire, 1998), with $C_{d}$ the drag coefficient. The rate of change of energy per unit area due to skin drag is proportional to $|u_{\text{orb}}-v|^{3}$ (Shen & Squire, 1998), where $u_{\text{orb}}=U\sin{(\phi)}$ is the wave orbital velocity at the ice-water interface, here evaluated at the free surface under the hypothesis of thin sea ice, and $v$ is the ice drift velocity. The magnitude of the orbital velocity is $U=a\Omega$, where $\Omega=gk/\omega$, and $\phi$ is time-dependent phase. Phase averaging, we obtain:
$s_{\text{sd}}=-\frac{C_{d}}{g}\frac{1}{2\pi}\int_{0}^{2\pi}\left|a\Omega\sin(\phi)-v\right|^{3}d\phi.$ (4)
@@ 2.1 Solving the Amplitude Profile
Equation 5 represents a challenging nonlinear ordinary differential equation. Here we aim to derive an analytical solution and to facilitate the mathematical treatment we assume the parameters $C_{d}$ and $\alpha$ to be spatially homogeneous (constant). The model can be extended for heterogeneous conditions but numerical solutions should be sought. A polynomial in $a$ for the integral of the skin drag (Eq. 4) can be found explicitly in closed form for $|v|\geq a\Omega$, see Appendix A. In effort to enable closed form analytical solutions we seek to an approximate polynomial form using an asymptotic expansion of the integrand for $|v|\ll a\Omega$ that are then extended into the region $|v|<a\Omega$ (see Appendix A). We obtain:
@@
Note that, in oceanic conditions, $c_{g}$ is two orders of magnitude lager than the wind and current induced drift, i.e. $\mathcal{O}(10)\,\mathrm{m/s}$ versus $\mathcal{O}(0.1)\,\mathrm{m/s}$, therefore $|v|\ll c_{g}$ implying that, with $\alpha\text{ and }\Gamma$ positive, Eq. 7 is strictly positive, therefore the amplitude $a$ decreases monotonically in Eq. 5. For $a_{0}\Omega>|v|$, we define $\hat{x}^{*}$ to be the location into the ice domain such that $a(\hat{x}^{*})\Omega=|v|$. Therefore, the condition $|v|\lessgtr a\Omega$ corresponds to $\hat{x}\lessgtr\hat{x}^{*}$ in the physical domain. In contrast, for $a_{0}\Omega\leq|v|$, one obtains $\hat{x}^{*}=0$.
@@
where $\mathbf{1}$ is the indicator function, equal to $1$ if the condition is true and 0 otherwise. The piecewise solution components result in:
@@
which controls the roots of the quadratic polynomial in $a$ affecting the solution in the $\hat{x}<\hat{x}^{*}$ region (see Appendix B). Note that the physical interpretation of $\delta$ will be discussed in §2.2.
The solution has a few interesting features. The amplitude profile is continuous in $\hat{x}$ (see Appendix B). The solution is nonlinear, i.e. drag effects and other mechanisms cannot be added as linear superposition. For $\hat{x}<\hat{x}^{*}$, the different branches of the solution in Eq. 9 differ in form but possess similar qualitative behaviour, where we note that $\hat{x}^{*}$ depends on $\delta$. For a wave with initial orbital velocity magnitude $U_{0}<|v|$ we obtain the solution in region B, $a_{B}$, applies to the entire domain $\hat{x}\geq 0$ .
@@
such that $a_{B}(\hat{x}=\hat{x}_{\text{end}})=0$.The extinction location is always in region B because it is where amplitude tends to zero and the condition $a\Omega<|v|$ is satisfied. The extinction location is unique to dissipation due to drift-induced skin drag and defines the extent of the wave affected region. It moves closer to the sea ice edge for increasing $C_{D}$ coefficient and $\alpha$ (see solution in red, for $v=0.05$ m/s and $\alpha=7\times 10^{-6}{\text{ m}^{-1}}$   $\delta=3.4$], vs in blue, for $v=0.05$ m/s and $\alpha=0$ $\delta=\infty$], in Fig. 1a, while its dependence on drift is not trivial. Such extinction locations are not present in the solutions of Kohout et al. (2011) and Herman et al. (2019), shown dashed green and orange respectively in Fig. 1a), for which $\alpha=0$ and $v=0$.
@@
Figure 1: Sample amplitude (a) and corresponding attenuation (b) profiles according to our model and formulations by Kohout et al. (2011) and Herman et al. (2019). Dashed lines denote $x^{*}$ which denotes the separation between region A and B. Amplitude is normalised by the initial amplitude and distance by $x_{\text{end}}$ for $\alpha=0$. Attenuation rate is normalised by $\alpha$. Solutions are shown for $a_{0}=1$ m, $C_{d}=0.05$, $v=0.053\,\mathrm{m\,s^{-1}}$, $\omega=0.63\,\mathrm{s^{-1}}$ , $\alpha=7.2\times 10^{-6}\mathrm{\,m^{-1}}$.
Fig. 1a displays the normalised amplitude as a function of the distance into the MIZ normalised by the value $x_{\text{end}}$ computed for the non zero drift and exponential attenuation case ($v=0.053\mathrm{m\,s^{-1}}$ and $\alpha=0$; the case displayed in blue). The attenuation is stronger when drift is present as shown in red ($v>0$ and $\alpha=7\times 10^{-6}$) compared to the solution in purple with ($v=0$ m/s and $\alpha=7\times 10^{-6}$). The increased attenuation due to other mechanisms also manifests when drift is present, see solutions in blue vs in red.
@@ 2.2 Predicting the Attenuation Rate
To allow better comparison with attenuation results for non-drifting sea ice as they appear in the literature (Kohout et al., 2011; Herman et al., 2019) and better highlight the attenuation characteristics,
we compute the effective amplitude attenuation rate in the moving frame in the ice covered region defined as
@@
with the global minimum located in the region $\hat{x}<\hat{x}^{*}$ if $a_{0}\Omega\geq 3\sqrt{2}|v|/2$, whereas for lower orbital velocity at the sea ice edge the attenuation monotonically increases. This result allows for a physical interpretation of the dimensionless parameter $\delta$: for $\delta>1$ wave attenuation is dominated by drift-induced skin drag and for $\delta<1$ all other mechanisms dominate.
For drifting sea ice, the attenuation rate grows unbounded in the neighbourhood of the extinction location (see solutions in red and blue in Fig. 1b), i.e.  $\alpha_{\text{eff}}(\hat{x})\sim[2(x_{\text{end}}-\hat{x})]^{-1}$, therefore the extinction location affects the magnitude of the attenuation rate in this region. When drift is absent, $\hat{x}^{*}\rightarrow\infty$ and the solution for $\hat{x}<\hat{x}^{*}$ applies to the entire domain. As a consequence of our model, skin drag dominates attenuation close to the sea ice edge whereas other processes dominate deeper into the sea ice, where we retrieve the classical exponential attenuation, i.e. $\alpha_{\text{eff}}\rightarrow\alpha$ for $\hat{x}\rightarrow\infty$ (see solution in purple in Fig. 1b; the grey dashed line denotes $\alpha_{\text{eff}}=\alpha$).